[Brahms-l] Shift report 20050619 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 19 2005 - 00:04:04 EDT
Shifter leader was SB
Shifters were 


1. Good beam the whole shift(beam extended till 2:30 AM)<br>

2. Runs 15623,15624, 15629-15636<br>

Trig2(766k), Trig3(244k), Trig4(139k) ,Trig5(993k)<br>
<br>
3. AC failure caused power in the FE Hut to shut off.<br>
Problem fixed and we are running for physics<br>
--------------------------------------<br>
Time   Blue Pol.(%)     Yellow Pol.(%)<br>
---------------------------------------<br>
16:30  -43.09 +/- 2.36  -51.52 +/- 2.21 <br>
19:30  -44.72 +/- 2.37  -41.54 +/- 2.30<br>
22:30  -44.99 +/- 2.33  -47.06 +/- 2.28<br>


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jun 19 00:04:20 2005

This archive was generated by hypermail 2.1.8 : Sun Jun 19 2005 - 00:04:27 EDT